OpenStudy (anonymous):

HELP AND U GET A MEDAL AND FAN If the graph of the function y=x^2-6 were shifted 3 units down what would the new shifted equation be

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

im coming to save you

OpenStudy (anonymous):

I think the new equation would be y= (x^2-6) -3

OpenStudy (anonymous):

what lol

OpenStudy (anonymous):

never mind i tried

OpenStudy (anonymous):

ok and that was not an choice

OpenStudy (anonymous):

well what are the choices it would make it easier to help if you provided us with the choices.

OpenStudy (anonymous):

oh hold on

OpenStudy (anonymous):

A. Y-3X^2-6 B. Y=X^2-3 C. Y=X^2-9 D. Y=3X^2-3

OpenStudy (anonymous):

HELLO

OpenStudy (anonymous):

I'm 99.9% positive it's A. but I could be wrong.. pretty sure I'm right though

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

It's D

OpenStudy (anonymous):

hope I helped :)

OpenStudy (anonymous):

wait is it A or D

OpenStudy (anonymous):

It's D